# DotResize: Reducing LLM Width via Discrete Optimal Transport-based Neuron Merging

### v1

@@ Abstract
Model compression offers a promising path to reducing the cost and inaccessibility of large pre-trained models, without significantly compromising their impressive performance. Large Transformer models, including large language models (LLMs), often contain computational redundancy, which can serve as a target for new model compression methods. In this work, we specifically target neuron-level redundancies in model layers by combining groups of similar neurons into fewer neurons. We frame this width reduction as a Discrete Optimal Transport problem, and propose DotResize, a novel Transformer compression method that uses optimal transport theory to transform and compress model weights. To ensure applicability within the Transformer architecture, we motivate and incorporate entropic regularization and matrix factorization into the transportation maps produced by our method. Unlike pruning-based approaches which discard neurons based on importance measures, DotResize  re-projects the entire neuron width, allowing the retention and redistribution of useful signal across the reduced layer. Empirical results show that compared to simple or state-of-the-art neuron width-pruning techniques, DotResize  can outperform these methods across multiple LLM families and sizes, while achieving measurable reductions in real-world computational cost.
@@ 1 Introduction
High-quality compression techniques for large language models (LLMs) are increasingly important due to the ubiquity of pre-training and success of model and data scaling. Compression can reduce model size and latency, which helps facilitate reduced inference cost, on-device capabilities, and improved access to models by both users and researchers. Although prior work has shown the theoretical success of weight pruning techniques (frantar2023sparsegpt, ; sunsimple, ), practical gains in on-disk compression and speed-up often depend on specialized hardware and specific tensor implementations. For example, unstructured and semi-structured pruning methods do not yield memory savings without dedicated sparse storage formats. Moreover, while semi-structured pruning can offer inference speed-ups, these benefits are limited to specific GPU architectures and typically come with greater performance degradation than unstructured approaches. Structured pruning techniques provides alternatives that solve some of these issues by design, where entire structures like attention heads, blocks of weight matrices, neurons, or entire layers can be removed, leading to direct inference speed-ups and memory savings. One variety of structured pruning involves reducing the width of weight matrices in model layers, resulting in compressed models with more computational throughput (ma2023llm, ; ashkboos2024slicegpt, ).
@@
Parallel to pruning-based methods, a complementary line of research has focused on direct parameter intervention that exploits functional invariance to change model parameters without changing the underlying model function. In other words, these interventions involve designing functions $\pi$ that act on model parameters $\theta$ such that the model preserves $f(x|\theta)=f(x|\pi(\theta))$. This functional invariance is central to many model merging methods (singh2020model, ; ainsworthgit, ), and has recently been explored for single-model compression applications. These applications tend to design functions $\pi^{*}$ such that the model function is approximated due to the compression objective, i.e. $f(x|\theta)\approx f(x|\pi^{*}(\theta))$. Examples include QuaRot, where rotational matrices are applied to Transformer weights to reduce outliers for improved low bit quantization (ashkboos2024quarot, ), and SliceGPT, where rotational PCA matrices are again applied to prepare weights for low-eigenvalue neuron pruning (ashkboos2024slicegpt, ).
In this work, we take an approach similar to structured pruning, building upon previous work that prunes the width of individual layers in LLMs ma2023llm ; ashkboos2024slicegpt . However, we employ a key difference: instead of determining neurons unimportant to model inference, we instead focus on combining neurons into fewer according to their activation similarities. In a direct parameter intervention approach without additional training, we focus on OT-based techniques to target redundant processing, a known phenomenon explored in previous work (dalvi-etal-2020-analyzing, ; nanda2023diffused, ). This approach is inspired by model merging methods where combining parameters from disparate models via arithmetic operations without additional gradient updates is enough to create high-quality multi-task models yang2024model . Prior work has examined the use of transportation maps that align neurons from different models’ layers singh2020model ; pena2023re ; imfeldtransformer , but this line of work assumes that there exists a clear alignment of neurons, as these models are generally trained on the same data, but from different initializations. We apply a similar principle, but without this guiding assumption in the challenging setting of parameter width reduction. We find that compared to both a simple activation pruning strategy and SliceGPT, adding OT-merging can further improve performance by incorporating additional signal from neurons that would otherwise be pruned. Our method can be summarized in Figure 1.
@@
Our contributions are as follows.
@@
    We propose a novel, training-free neuron width reduction method that uses discrete optimal transport theory to frame the compression problem, in order to compress signal from the full neuron width rather than discard signal from pruned neurons.
@@
    We examine the computational invariance of Transformer weights to orthogonal maps proposed in Ashkboos et al. ashkboos2024slicegpt , and extend the set of invariant maps from orthogonal to any invertible map, via the application of QR-decomposition to respect the orthogonal-only invariance of RMSNorm.
@@ 2 Related Work
Given the practical limitations of unstructured weight pruning (lecun1989optimal, ), including the lack of real-world memory and latency savings, prior work has examined converting unstructured sparsity patterns into semi-structured sparsity patterns (frantar2023sparsegpt, ; sunsimple, ). However, this conversion process leads to additional performance loss and only provides speed-ups on certain GPU architectures. Alternatively, structured pruning techniques that remove structured portions of models, like attention heads, layers, individual neurons, and blocked weight regions, provide a hardware agnostic approach to achieving both real-world disk and latency savings. In this work, we focus on reducing individual neurons, which has been explored in prior work via neuron width pruning. Ma et al. ma2023llm propose a “channel pruning” technique that removes a neurons sharing the same index across all LLM layers, effectively reducing the width of the entire model. Ashkboos et al. ashkboos2024slicegpt propose a related technique, but use PCA projections to perform the width reduction to prune low eigenvalue associated neurons in the re-projected eigenspace. Gao et al. gao2024disp reduce neuron width in a dimension-independent manner by using different indexing matrices per layer. However, their method trains a separate GRU hypernetwork to predict neurons to remove. We depart from these pruning approach by instead focusing on merging-based strategies.
@@
Current LLMs are highly overparameterized, containing more parameters than necessary to fit their training data. This leads to consequences like the redundant encoding of concepts, which has been repeatedly identified in prior work. Srinivas and Babu srinivas2015data propose a method to wire similar neurons together within a layer based on their weight set similarities. Dalvi et al. dalvi-etal-2020-analyzing analyze redundancy between individual neurons and layers and propose a clustering-based approach to pruning neurons. In a model merging method, Stoica et al. stoicazipit tackle redundancy by not only aligning similar neurons across models to merge, but also align similar neurons within a model for merging. Nanda et al. nanda2023diffused characterize redundancy in pre-trained models as “diffuse” where multiple random subsets of neurons in a model layer can approximate the performance of the whole layer. In our work, we are inspired by this diffusely redundant nature of neurons, and design our technique to target potentially redundant groups of neurons within a layer.
@@ 3 DotResize
Our DotResize  width reduction method draws on entropy-regularized discrete optimal transport, where we assign a layer’s neurons to a source distribution and a subset of these neurons as a target distribution, and compute an optimal transport map between source and target based on activation signatures over representative data. The result of this approach is a method that reduces overall model width by transforming the signal from the entire neuron width, which is in direct contrast with pruning methods that only consider a portion of the signal. A demonstration of this width reduction can be seen in Figure 1. Reducing model width incidentally reduces activation width too, improving activation memory use.
@@ 3.1 Computing the transport maps
Discrete optimal transport problems are formulated as transforming a source distribution to a target distribution subject to a cost matrix that dictates the cost to transform each source item to each target item. Following Singh et al. singh2020model , we define our source and target distributions as uniform, with the source as $\mathbf{a}=\mathbf{1}_{d_{\text{orig}}}/d_{\text{orig}}$, and target as $\mathbf{b}=\mathbf{1}_{d_{\text{new}}}/d_{\text{new}}$, where we select $d_{\text{new}}<d_{\text{orig}}$ neurons from the original width. While selection strategies can differ, we select using average activation norm in our work, as detailed in later sections. We define the cost matrix of transporting mass from distribution $\mathbf{a}$ to distribution $\mathbf{b}$, also known as ground metric, as the pairwise distances between the activations over $d_{\text{orig}}$ source neurons, and the subset of activations corresponding to the $d_{\text{new}}$ selected neurons. This means that at each layer we wish to compress, we compute activations across $t$ exemplar tokens and use these activations to help determine our ground metric. In our work, we use $\ell_{1}$ norm to compute pairwise distances. This results in cost matrix $C\in\mathbb{R}^{n\times k}$, where for rows of activations $x_{i},x_{j}\in\mathbb{R}^{t}$, $C_{ij}=||x_{i}-x_{j}||_{1}$.
In computing the discrete optimal transport map that dictates how to recombine these $d_{\text{orig}}$ neurons into $d_{\text{new}}$, we incorporate Sinkhorn entropic regularization into the OT optimization problem in order to incorporate “soft” alignments between neurons (cuturi2013sinkhorn, ). “Soft” alignments allow for each neuron to correspond to linear combinations of other neurons, whereas non-regularized alignments are much sparser. Additionally, Sinkhorn solutions are less costly to compute compared to non-regularized optimal transport distances. The following optimization problem reflects the Sinkhorn objective for finding map $T$ subject to costs $C\in\mathbb{R}^{n\times k}$.
$T=\arg\min_{T}\langle T,C\rangle-\lambda H(T)$ (1)
The larger the parameter $\lambda$, the “softer” the transportation map is due to the entropic regularization reflected in $H(T)$. In practice, we multiply solution $T$ by $d_{\text{orig}}$ so that each row sums to 1 rather than $\frac{1}{d_{\text{orig}}}$ as is the case in optimal transport solutions. After applying map $T$ to the one or more weight matrices that produced the activations used for alignment, we can apply an inverse transformation to the following weight matrix such that functional invariance is preserved. In practice, we must modify these maps in order to work with the Transformer architecture, as described in following sections.
@@ 3.2 Maintaining invariance via QR decomposition
Figure 2: Our QR-decomposition step allows for general invertible matrices to apply at residual junctions as depicted by the figure. The figure depicts a general residual connection as $\bigoplus$ in a pre-norm Transformer layer. While orthogonal transformations are naturally invariant to RMSNorm, general invertible matrices are not unless decomposed via QR decomposition. The commutativity of matrix multiplication allows us to absorb matrix R into the $T^{-1}$ calculation, allowing the orthogonal multiplicand to not change RMSNorm.
In applying the transportation matrices to transform model weights, we aim to approximate the original function to achieve a compressed model without losing substantial performance. In other words, our transportation maps, when applied to weights, should not substantially change the overall function of the model. In Ashkboos et al. ashkboos2024slicegpt , orthogonal PCA matrices are used to transform model weights given their invariance to the commonly used root mean square layer normalization (RMSNorm) in Transformers (zhang2019root, ). This invariance allows for transformation matrices to apply to weights before and after a residual connection, while not interfering with the function of RMSNorm in pre-norm architectures. For orthogonal matrix $Q$ and data vector $\mathbf{x}$, the relationship can be expressed as the following:
@@
A depiction of this invariance is found in Figure 2. This invariance assumes that RMSNorm is unweighted, as weighted versions can be converted to unweighted RMSNorms by pre-multiplying weight matrices following pre-norm by the RMSNorm weights. For more details on this conversion, refer to Ashkboos et al. ashkboos2024slicegpt .
To apply our method, we relax the constraint of permissible transportation matrices from only orthogonal matrices ($Q$) to any invertible matrix. However, the original use of orthogonal matrices in prior work was important as they do not change vector norms, whereas general invertible matrices may. To address this limitation of invertible transformation matrices, we use QR decomposition in order to decompose invertible transformation matrix $T=QR$ such that the $Q$ matrix is applied to the input vector before RMSNorm, and the $R$ matrix is applied after the RMSNorm. We maintain the unweighted RMSNorm and pre-norm assumptions, and convert weighted RMSNorms to unweighted RMSNorms as previously described. Given decomposition $T=QR$, inverse matrix $T^{-1}$ and data vector $\mathbf{x}$, we have the following relationship:
$\text{RMSNorm}(\mathbf{x}Q)RT^{-1}=\frac{\mathbf{x}Q}{\lVert\mathbf{x}Q\rVert}RT^{-1}=\frac{\mathbf{x}QRT^{-1}}{\lVert\mathbf{x}\rVert}=\text{RMSNorm}(\mathbf{x})$ (3)
A simple explanation of steps in this equality appears in Appendix A. Given the associativity of matrix multiplication, we are able to QR-decompose our computed transport map and then store just $M=Q$ and $M^{\text{inv}}=RT^{-1}$ as our transport and untransport maps. An overview of the application of these maps is summarized in Figure 2.
While in the case of full rank $T$ this equality holds, we note that our transport maps are not full rank due to the intentional difference in the width of source and target supports. In our setting, the orthogonal matrix resulting from QR-decomposition is still critical to the application of our method, so much so that our method does not converge without its application due to its role in preserving layer normalization. Additionally, we use the pseudoinverse for non-invertible $M$, and find that this approximation also works well.
@@ 3.3 Applying maps
With this QR modification, we are able to apply our transformations at two key locations for each Transformer layer vaswani2017attention . We label our transformations per layer as $\{M_{A},M_{A}^{\text{inv}}\}$ and $\{M_{F},M_{F}^{\text{inv}}\}$.
@@
  • $W_{Q},W_{K},W_{V}$ are the $d_{\text{orig}}\times d_{\text{orig}}$ weight matrices for queries, keys, and values in multi-headed (or grouped-query ainslie2023gqa , where dimensions are $d_{\text{orig}}\times(d_{\text{orig}}/n_{\text{groups}})$) attention
@@
  • $W_{\text{up}}$ and $W_{\text{gate}}$ are the $d_{\text{orig}}\times d_{\text{ff}}$ up-projection matrices in feed-forward layers with GLU-based activations, which we denote as $\sigma$ for the nonlinearity and $\odot$ for gating shazeer2020glu .
@@
$x_{\text{out}}^{\text{attn}}=W_{O}\text{MHA}(W_{Q}M_{F}^{\text{inv}}x_{\text{in}}^{\text{attn}\prime},W_{K}M_{F}^{\text{inv}}x_{\text{in}}^{\text{attn}\prime},W_{V}M_{F}^{\text{inv}}x_{\text{in}}^{\text{attn}\prime})+M_{F}^{\text{inv}}x_{\text{in}}^{\text{attn}\prime}$ (4)
@@
$M_{A}x_{\text{out}}^{\text{attn}}=M_{A}W_{O}\text{MHA}(W_{Q}M_{F}^{\text{inv}}x_{\text{in}}^{\text{attn}\prime},W_{K}M_{F}^{\text{inv}}x_{\text{in}}^{\text{attn}\prime},W_{V}M_{F}^{\text{inv}}x_{\text{in}}^{\text{attn}\prime})+M_{A}M_{F}^{\text{inv}}x_{\text{in}}^{\text{attn}\prime}$ (5)
Now, to reverse the effect of $M_{A}$ on $x_{\text{out}}^{\text{attn}}$, which we will denote as $x_{\text{in}}^{\text{ff}\prime}=M_{A}x_{\text{out}}^{\text{attn}}$, we have:
$x_{\text{out}}^{\text{ff}}=W_{\text{down}}\sigma\left((W_{\text{up}}M_{A}^{\text{inv}}x_{\text{in}}^{\text{ff}\prime})\odot(W_{\text{gate}}M_{A}^{\text{inv}}x_{\text{in}}^{\text{ff}\prime})\right)+M_{A}^{\text{inv}}x_{\text{in}}^{\text{ff}\prime}$ (6)
@@
$M_{F}x_{\text{out}}^{\text{ff}}=M_{F}W_{\text{down}}\sigma\left(W_{\text{up}}M_{A}^{\text{inv}}x_{\text{in}}^{\text{ff}\prime}\odot W_{\text{gate}}M_{A}^{\text{inv}}x_{\text{in}}^{\text{ff}\prime}\right)+M_{F}M_{A}^{\text{inv}}x_{\text{in}}^{\text{ff}\prime}$ (7)
All $M$ matrices can be “absorbed” into neighboring weight matrices $W$ due to the associativity of matrix multiplication. This application of maps mirrors several prior methods using a similar project/unproject scheme for Transformer model merging and compression imfeldtransformer ; ashkboos2024slicegpt ; verma2024merging .

### v2

@@ Abstract
Structured pruning methods designed for Large Language Models (LLMs) generally focus on identifying and removing the least important components to optimize model size. However, in this work, we question this prevalent approach by instead exploring how to recombine information from structures designated for pruning back into the reduced model. We specifically focus on neuron width reduction, and frame this problem as a Discrete Optimal Transport problem, and propose DotResize, a novel Transformer compression method that uses optimal transport theory to transform and compress model width. To ensure applicability within the Transformer architecture, we motivate and incorporate necessary entropic regularization and matrix factorization techniques into the transportation maps produced by our method. Unlike pruning-based approaches which discard neurons based on importance measures, DotResize  re-projects the entire neuron width, allowing the retention and redistribution of useful signal across the reduced layer. Empirical results show that compared to simple or state-of-the-art neuron width-pruning techniques, DotResize  serves as a useful add-on to pruning, while achieving measurable reductions in real-world computational cost.
Optimal Transport, LLM Pruning, Merging
@@ 1 Introduction
High-quality compression techniques for LLMs have become increasingly important due to the ubiquity of pre-training and success of model and data scaling. Compression can reduce model size and latency, which helps facilitate reduced inference costs, on-device capabilities, and improved access to models by both users and researchers.
Structured pruning is a popular approach to compression because it directly reduces the on-disk size of LLMs and accelerates inference, without requiring specialized hardware or libraries, often needed for unstructured pruning. By targeting and removing entire structures, such as attention heads, blocks from weights, neurons, or entire layers, structured pruning delivers direct inference speed-ups and memory savings. One variety of structured pruning reduces the width of weights across model layers, resulting in compressed models with more computational throughput (Ma et al., 2023; Ashkboos et al., 2024a). In general, these width pruning techniques aim to identify important neurons, and discard neurons which are deemed less important.
@@
Beyond weight removal, a complementary line of research focuses on applying invariant transformations to model parameters. These methods involve designing functions $\pi$ that modify model parameters $\theta$ such that the model preserves $f(x|\theta)=f(x|\pi(\theta))$, to prepare the model for downstream tasks. Prior work has designed invariant transformations to enable many model merging methods (Singh & Jaggi, 2020; Ainsworth et al., 2023), and has recently been explored for single-model compression applications (Theus et al., 2024). These single-model applications tend to design functions $\pi^{*}$ such that the model function is approximated due to the compression objective, i.e. $f(x|\theta)\approx f(x|\pi^{*}(\theta))$. Relevant examples include QuaRot, where rotational matrices are applied to Transformer weights to reduce outliers for improved low bit quantization (Ashkboos et al., 2024b), and SliceGPT, where rotational PCA matrices are again applied to prepare weights for low-eigenvalue neuron pruning (Ashkboos et al., 2024a).
@@
In this work, we move beyond the binary choice of keeping or discarding neurons inherent to pruning by proposing an merging-based augmentation to the common neuron pruning paradigm (Ma et al., 2023; Ashkboos et al., 2024a). We introduce a method that merges signal from neurons to be pruned into the remaining width using optimal transport. This approach represents a first step toward unifying model merging and structured pruning, aiming to reincorporate the “unaccounted” signal that is typically lost during compression. While a prior transport-based approach has been proposed and applied to ResNet-based image models Theus et al. (2024), we identify and propose solutions to overcome the key architectural hurdles required to apply these techniques to LLMs. Our method does not require recovery fine-tuning, offering a promising new direction for efficient, zero-shot model compression.
We find that when applied atop of gradient-free LLM neuron pruning methods, DotResize frequently recovers additional performance by re-incorporating signal from neurons that would have otherwise be pruned. Our method is summarized in Figure 1.111Full code including an implementation of our method and experiments will be released upon publication.
Our contributions can be summarized as follows.
@@
    We propose a novel, training-free neuron width reduction method in LLMs that uses entropy-regularized discrete optimal transport to recombine signal from neurons that would otherwise be pruned from neuron pruning methods.
@@
    We re-examine a specific invariance result found in Ashkboos et al. (2024a) that enables the application of orthogonal maps throughout LLM weights. We extend this set of invariant maps from orthogonal to any invertible map, via the application of QR-decomposition to respect the orthogonal-only invariance of RMSNorm.
@@ 2 Related Work
Given the practical limitations of unstructured weight pruning (LeCun et al., 1989), including the lack of easily realizable real-world memory and latency savings, recent work on weight pruning has examined converting unstructured sparsity patterns into semi-structured sparsity patterns (Frantar & Alistarh, 2023; Sun et al., 2024). However, this conversion process frequently leads to additional performance loss and only provides speed-ups on specific GPU architectures. Alternatively, structured pruning techniques that remove structured portions of models, like attention heads, layers, individual neurons, and blocked weight regions, provide a hardware agnostic approach to achieving both real-world disk and latency savings. In this work, we focus on reducing individual neurons, which has been explored in prior work via neuron width pruning. Ma et al. (2023) propose a “channel pruning” technique that removes neurons sharing the same index across all LLM layers, effectively reducing the width of the entire model. Ashkboos et al. (2024a) propose a related technique, but use PCA projections to perform the width reduction to prune low eigenvalue associated neurons in the re-projected eigenspace. Gao et al. (2024b) reduce neuron width in a dimension-independent manner by using different indexing matrices per layer. Their method trains a separate GRU hypernetwork to predict neurons to remove. We depart from these pruning approaches by instead focusing on merging-based strategies.
@@
Current LLMs are highly overparameterized, containing more parameters than necessary to fit their training data. This leads to consequences like the redundant encoding of concepts, which has been repeatedly identified in prior work. Srinivas & Babu (2015) propose a method to wire similar neurons together within simple networks based on their weight set similarities. Dalvi et al. (2020) analyze redundancy between individual neurons and layers and propose a clustering-based approach to pruning neurons. In a model merging method, Stoica et al. (2024) tackle redundancy by not only aligning similar neurons across models to merge, but also align similar neurons within a model for merging. Nanda et al. (2023) characterize redundancy in pre-trained models as “diffuse,” where multiple random subsets of neurons in a model layer can approximate the performance of the whole layer. In our work, we are inspired by this diffusely redundant nature of neurons, and design our technique to target potentially redundant groups of neurons within a layer. Theus et al. (2024) use optimal transport to merge groups of neurons specifically in ResNets, where only certain residual groups undergo optimal-transport based compression. We extend ideas in this approach to fully residually-connected LLMs, and operate at a layer level rather than group-level.
@@ 3 DotResize
Our DotResize width reduction method draws on entropy-regularized discrete optimal transport, where we assign a given layer’s neurons to a source distribution and a subset of these neurons as a target distribution, and compute an optimal transport map between source and target based on activation signatures over representative data. The result of this approach is a method that reduces overall model width by transforming the signal from the entire neuron width, which is in contrast with pruning methods that only consider a portion of the signal. A demonstration of this width reduction is displayed in Figure 1. Reducing model width incidentally reduces activation width too, improving activation memory use.
@@ 3.1 Computing the transport maps
Discrete optimal transport problems are formulated as transforming a source distribution $\mathbf{a}$ to a target distribution $\mathbf{b}$ subject to a cost matrix that dictates the cost of transforming each source item to each target item, denoted as $C$. The goal of optimal transport is to find a map that minimizes this cost.
[…]
We find our transportation maps via an entropy-regularized objective (Cuturi, 2013). We use Sinkhorn regularization to encourage smoothing of the transportation map, such that each target neuron can be a combination of multiple source neuron, rather than just a few. Prior work has found Sinkhorn regularization to be critical when merging separate Transformers via transport mapping (Imfeld et al., 2024). Additionally, Sinkhorn regularization improves the computational tractability of the OT objective.
The following optimization problem reflects the Sinkhorn objective for finding map $T$ between distributions $\mathbf{a}$ and $\mathbf{b}$ subject to costs $C$.
$T=$ $\arg\min_{T}\langle T,C\rangle-\lambda H(T)$
$\text{s. t. }T\mathbf{1}_{d_{\text{orig}}}=\mathbf{a},T^{T}\mathbf{1}_{d_{\text{new}}}=\mathbf{b}$ (1)
The larger the parameter $\lambda$, the “softer” the transportation map is due to the flattening effect of the entropy term.
After applying map $T$ to the one or more weight matrices that produced the activations used for alignment, we can apply an inverse transformation to the following weight matrix such that functional invariance is preserved. In practice, we must modify these maps in order to interface with the Transformer architecture, as described in following sections.
@@ 3.2 Maintaining invariance via QR decomposition
In applying the transportation matrices to transform model weights, we aim to approximate the original model’s function to achieve compression without losing substantial performance. In Ashkboos et al. (2024a), orthogonal PCA matrices are used to transform model weights given their invariance to the commonly used root mean square layer normalization (RMSNorm) in Transformers (Zhang & Sennrich, 2019). This invariance allows for transformation matrices to apply to weights before and after a residual connection, while not interfering with the function of RMSNorm in pre-norm architectures. For orthogonal matrix $Q$ and data vector $\mathbf{x}$, this invariance can be expressed as the following:
@@
A depiction of this invariance is found in the first panel of Figure 2. This invariance assumes that RMSNorm is unweighted, as weighted versions can be converted to unweighted RMSNorms by pre-multiplying weight matrices following pre-norm by the RMSNorm weights (Elhage et al., 2023). For more details on this conversion, refer to Ashkboos et al. (2024a).
Figure 2: Our QR-decomposition step allows for general invertible matrices to apply at residual junctions as depicted by the figure. The figure depicts a general residual connection as $\bigoplus$ in a pre-norm Transformer layer. While orthogonal transformations are naturally invariant to RMSNorm, general invertible matrices are not unless decomposed via QR decomposition and re-routed as shown. The associativity of matrix multiplication allows us to absorb matrix R into the $T^{-1}$ calculation, allowing the orthogonal multiplicand to not change RMSNorm.
To apply our method, we must relax the constraint of permissible transportation matrices from only orthogonal matrices to any invertible matrix. While the use of orthogonal matrices in prior work was important as they do not change vector norms, general invertible matrices may change norms. To address this limitation of invertible transformations, we use a QR decomposition step in order to decompose invertible matrix $T=QR$ such that the $Q$ matrix is applied to the input vector before RMSNorm, and the $R$ matrix is applied after the RMSNorm. Given decomposition $T=QR$, inverse matrix $T^{-1}$ and data vector $\mathbf{x}$, we have the following relationship:
$\text{RMSNorm}(\mathbf{x}Q)RT^{-1}=\frac{\mathbf{x}Q}{\lVert\mathbf{x}Q\rVert}RT^{-1}$
$=\frac{\mathbf{x}QRT^{-1}}{\lVert\mathbf{x}\rVert}=\text{RMSNorm}(\mathbf{x})$ (3)
A simple explanation of steps in this equality appears in Appendix A. Given the associativity of matrix multiplication, we are able to QR-decompose our computed transport map and then store just $M=Q$ and $M^{\text{inv}}=RT^{-1}$ as our final transport and untransport maps. An overview of the application of these maps is summarized in Figure 2.
While in the case of square $T$ this equality holds, we note that our transport maps are rectangular due to the intentional difference in the width of source and target supports. Rectangular $T$ is not invertible, meaning that we instead use the pseudoinverse instead in our method. We show that up to orthogonal projection onto $\text{col}(Q)$ this equality holds in Appendix B, and in practice, this QR modification is vital for stability; without it, our method fails to converge as the preservation of layer normalization is lost. Despite our specific application of these invariant maps in our setting, we believe this technique to be broadly useful in other methods that transform models in-place using invariant or approximately invariant maps.
@@ 3.3 Applying maps
With this QR modification, we are able to apply our transformations at two key locations for each Transformer layer (Vaswani et al., 2017). We label our transformations per layer as $\{M_{A},M_{A}^{\text{inv}}\}$ and $\{M_{F},M_{F}^{\text{inv}}\}$.
@@
  • $W_{Q},W_{K},W_{V}$ are the $d_{\text{orig}}\times d_{\text{orig}}$ weight matrices for queries, keys, and values in multi-headed attention (or grouped-query attention (Ainslie et al., 2023), where dimensions are $d_{\text{orig}}\times(d_{\text{orig}}/n_{\text{groups}})$)
@@
  • $W_{\text{up}}$ and $W_{\text{gate}}$ are the $d_{\text{orig}}\times d_{\text{ff}}$ up-projection matrices in feed-forward layers with GLU-based activations, which we denote as $\sigma$ for the nonlinearity and $\odot$ for gating (Shazeer, 2020).
@@
$x_{\text{out}}^{\text{attn}}=W_{O}\text{MHA}($ $W_{Q}M_{F}^{\text{inv}}x_{\text{in}}^{\text{attn}\prime},$
$W_{K}M_{F}^{\text{inv}}x_{\text{in}}^{\text{attn}\prime},$
$W_{V}M_{F}^{\text{inv}}x_{\text{in}}^{\text{attn}\prime})+M_{F}^{\text{inv}}x_{\text{in}}^{\text{attn}\prime}$ (4)
@@
$M_{A}x_{\text{out}}^{\text{attn}}=M_{A}W_{O}\text{MHA}($ $W_{Q}M_{F}^{\text{inv}}x_{\text{in}}^{\text{attn}\prime},$
$W_{K}M_{F}^{\text{inv}}x_{\text{in}}^{\text{attn}\prime},$
$W_{V}M_{F}^{\text{inv}}x_{\text{in}}^{\text{attn}\prime})+M_{A}M_{F}^{\text{inv}}x_{\text{in}}^{\text{attn}\prime}$ (5)
Now, to reverse the effect of $M_{A}$ on $x_{\text{out}}^{\text{attn}}$, which we will denote as $x_{\text{in}}^{\text{ff}\prime}=M_{A}x_{\text{out}}^{\text{attn}}$, we multiply our ff input by $M_{A}^{\text{inv}}$:
$x_{\text{out}}^{\text{ff}}=W_{\text{down}}\sigma$ $(W_{\text{up}}M_{A}^{\text{inv}}x_{\text{in}}^{\text{ff}\prime}$
$\odot W_{\text{gate}}M_{A}^{\text{inv}}x_{\text{in}}^{\text{ff}\prime})+M_{A}^{\text{inv}}x_{\text{in}}^{\text{ff}\prime}$ (6)
@@
$M_{F}x_{\text{out}}^{\text{ff}}=M_{F}W_{\text{down}}\sigma$ $(W_{\text{up}}M_{A}^{\text{inv}}x_{\text{in}}^{\text{ff}\prime}$
$\odot W_{\text{gate}}M_{A}^{\text{inv}}x_{\text{in}}^{\text{ff}\prime})+M_{F}M_{A}^{\text{inv}}x_{\text{in}}^{\text{ff}\prime}$ (7)
All $M$ matrices can be “absorbed” into neighboring weight matrices $W$ due to the associativity of matrix multiplication. This application of maps mirrors several prior methods using a similar project/unproject scheme for Transformer model merging and compression (Imfeld et al., 2024; Ashkboos et al., 2024a; Verma & Elbayad, 2024).